OpenStudy (anonymous):

Liz works two part time jobs for a total of 34 hours a week. Job A pays $6.00 per hour and Job B pays her $7.20 per hour. How many hours did she work at each job the week that she made $226.80?

OpenStudy (anonymous):

I thought I posted the answer... a+b=34 6a+7.2b=226.80 solve it & you'll get a=hours @ jobA & b=hours @job B problem solving it...?

OpenStudy (anonymous):

you need to break this down into two equations. lets look at it like this: lets say that she works x hours for job A and y hours for job B. we know that together she works 34 hours a week. so putting this in mathematical terms we get: x+y=34 now since we have two variables, we need two equations. our second equation has to do with the money, since we know how much money she makes at each job, we can simply multiply. x(6.00)+y(7.20)=226.80 now we need to find out what we dont know. we need to solve our first equation for either x or y and then we can plug it in to the second equation. lets solve it for x x=34-y plugging that in we get: (34-y)(6)+(y)(7.2)=226.8 204-6y+7.2y=226.8 7.2y-6y=226.8-204 1.2y=22 y=18.3333333333 so she worked 18 and 1/3 hours at job B. now lets plug that back into our original equation to get how many hours she worked at job A: x=34-(18.3333) x=15.6666666666 so she worked 15 and 2/3 hours at job A

OpenStudy (anonymous):

a little mistake... 1.2b=22.8 b=19 hours a=15 hours